OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of the function.

OpenStudy (anonymous):

\[\left(\begin{matrix}\lim \\ x->10-\end{matrix}\right)\frac{ 1 }{ x-10 }\]

OpenStudy (anonymous):

that would be x and arrow pointing to the 10 with a negative notation

OpenStudy (phi):

yes, it means the limit as x approaches 10 from the left side (less than 10) in other words, let x be 9.9 , 9.99, 9.999 what do you get for 1/(x-10) for each of those numbers?

OpenStudy (anonymous):

Well for the first I got -10

OpenStudy (phi):

you could do it with paper and pencil, or use a calculator, or use google 1/(9.9-10)= in google gives -10 now as x gets closer to 10, say x=9.99 what do you get ?

OpenStudy (anonymous):

-100

OpenStudy (phi):

yes, and x=9.999 gives -1000 if you plotted these pairs of numbers: (9.9, -10) (9.99,-100) (9.999,-1000) what do you see ?

OpenStudy (anonymous):

Oh okay they just add a 0

OpenStudy (phi):